# hibayes.bib

### v1

@@ 2.1 Multilevel (hierarchical) linear modeling using Generalized Linear Models
$\begin{split}Y_{i}&\sim\text{Binomial}(n_{i},p_{i})\\ \text{logit}(p_{i})&=\beta_{0}+\beta_{1}X_{1i}+\cdots+\beta_{k}X_{ki}\end{split}$ (1)
$Y_{i}$ denotes the binary outcome (0 or 1) of sample $i$, distributed as the proportion of successes (proportion $p_{i}$ in $n_{i}$ Bernoulli trials). The success probability $p_{i}$ is linked to the linear predictor via the logit link function.
@@ 3.1 Practical considerations for fitting GLMs using HiBayES
    • The number of effective samples (n_eff) is the number of samples that were drawn for the respective parameter. This number gives an indication of how well-defined the GLM is and how smoothly the sampling process proceeded. This should be $\geq$ the number of MCMC samples drawn per chain (4000).
@@
Figure 3: Highest posterior density intervals (HPDIs) used for hypothesis testing. Depicted are posterior probability distributions estimating two LLMs’ average performance and 95% HPDIs of these posterior distributions (both on probability scale) from simulated data fitted with a GLM. The HPDIs represent the interval of parameter values that is most consistent with the data – the range of parameter values that the true performance score is most likely to fall into. HPDIs in panel A overlap completely (>99% overlap), which suggests that LLM performance is equivalent (LLM A == LLM B). In panel B, HPDIs overlap partially ( 88% overlap), hence the result is inconclusive (LLM B <> LLM A). In panel C, HPDIs do not overlap (0% overlap), so we can conclude that LLMs perform differently (LLM B > LLM A).
@@ 3.3 Use case 1: One LLM evaluated on multiple domains (benchmarks) with varying numbers of items
Figure 5: Forest plot to display posterior means, posterior distributions, and 95% HPDI (on the probability scale) for use case 1. Focusing on domain-specific and overall performance effects (blue), this plot provides an indication of the width (thin lines) and uncertainty of the posterior parameter distribution (thick lines, 95% HPDI), around the posterior mean (dots) for the respective parameter estimates. Additionally, for comparison between HiBayES and conventional evaluation statistics approaches, we present empirical means and SEM per domain (orange) and for overall (red).
While the HiBayES GLM estimated domain-specific effects and the empirical means per domain are similar, the empirical SEMs are much narrower than the GLM estimated HPDIs. This may be misleading and produce overly confident interpretations of LLM capabilities, as the empirical SEM provides too narrow estimates of uncertainty. However, an even larger risk of overstated capability interpretations stems from the overestimation of overall performance levels using the empirical mean. Here, the estimate is influenced strongly by the larger number of datapoints from domain 2 ($N=6768$) vs domain 1 ($N=1257$). While this could potentially be mitigated by averaging per domain first and then computing the overall performance, such an approach will not be reliable for estimating the SEM. Using HiBayES the overall performance is estimated exactly at the central point between Coding and Reasoning estimates (as one would intuitively expect) and the HPDI spans the width of both domain-specific posterior distributions – the overall mean estimate regularizes the domain-specific estimates.
@@
$\text{logit}(p_{i,j,k})$ $=\text{model}_{i}+\text{domain}_{i,j}+\gamma_{\text{subdomain}_{i,j,k}}$
$\text{model}_{i}$ $\sim\text{Normal}(\bar{\mu}_{\text{model}},\sigma_{\text{model}})$
$\sigma_{\text{model}}$ $\sim\text{HalfNormal}(0,0.1)$
$\bar{\mu}_{\text{model}}$ $=\mu_{\text{overall}}+\sigma_{\text{overall}}\cdot z_{\text{overall}}$
@@
$\sigma_{\text{overall}}$ $\sim\text{HalfNormal}(1)$
$z_{\text{overall}}$ $\sim\text{Normal}(0,1)$
$\text{domain}_{i,j}$ $\sim\text{Normal}(\text{model}_{i},\sigma_{\text{domain}})$
@@
$\gamma_{\text{subdomain}_{i,j,k}}$ $\sim\text{Normal}(\text{model}_{i},\sigma_{\text{subdomain}})$
$\sigma_{\text{subdomain}}$ $\sim\text{HalfNormal}(0.1).$
@@
$\text{Correct}_{i,j,k}$ denotes the number of correct answers for the $i$-th LLM, $j$-th domain, and $k$-th subdomain, distributed as the proportion of correct answers (proportion parameter $p$) among the number of trials ($n_{i,j,k}$). The GLM assumes that the individual LLM-specific parameters are drawn from an LLM distribution, specified as a Normal distribution. The individual domain-specific parameters are drawn from a domain distribution that is parameterized by the LLM-specific effects. The individual subdomain parameters are drawn from a subdomain distribution that is parameterized by the domain-specific effects.
@@
Figure 6 shows the posterior distributions and compares estimated model (LLM), domain and subdomain effects and the uncertainty of the estimates (HPDI) to empirical means (arithmetic means) and standard errors of the mean (SEM) for each LLM, domain and subdomain separately (Figure 6). Importantly, the empirical means per LLM, domain and subdomain are generally higher than the GLM estimated posterior means. The empirical SEMs are substantially narrower than the GLM estimated HPDIs. This phenomenon is called “shrinkage” and relates to hierarchical GLMs partially pooling parameter estimates across the nested levels of the data hierarchy – this form of regularization imposed by the hierarchical data structure produces less over-fitting and overconfident estimates. In other words, the parameter estimates derived from hierarchical GLMs allow us to more confidently generalize the capability in question to new contexts beyond the specific evaluation – e.g., real-world deployment contexts.
@@
Figure 6: Forest plot to display posterior means, posterior distributions, and 95% HPDI (on the probability scale) for use case 2. Focusing on model (LLM)-, domain- and subdomain-specific effects, this plot provides an indication of the width (thin lines) and uncertainty of the posterior parameter distribution (thick lines, 95% HPDI), around the posterior mean (blue dots) for the respective parameter estimates. Additionally, for comparison between HiBayES and conventional evaluation statistics approaches, we present empirical means and SEM per LLM, domain, subdomain (orange).
The HPDIs for the subdomains estimated in a hierarchical GLM do not suggest differences found using the conventional approach (95%-HPDI DS-1000 = [.10;.43].10.43[.10;.43][ .10 ; .43 ] vs [.13;.52].13.52[.13;.52][ .13 ; .52 ] and 95%-HPDI BoolQ = [.33;.77].33.77[.33;.77][ .33 ; .77 ] vs [.36;.80].36.80[.36;.80][ .36 ; .80 ]). Conventional statistics result in a large and statistically significant difference in LLM performance which is not supported by hierarchical GLMs that adequately reflect the data structure. Using HiBayES, statistical inferences and the related claims are more balanced, nuanced, and not primarily driven by small variance due to large evaluation sample sizes.
@@
An additional advantage of HiBayES is that GLMs like the one shown in Equation 3 straightforwardly extend to evaluation scenarios that cover more than two domains or more than two LLMs (e.g., comparing 3 LLMs) without adjusting the statistical inference process, correcting for multiple comparisons, or using another family of tests – we illustrate this below in use case 3. Since all estimated parameters are indexed, it is possible to add another number/index and estimate 3 (x3) LLM/domain parameters instead of 2 (x2).
@@
Interest in how AI models behave in real-world, agentic settings has surged in recent years. To test these capabilities, challenging and extended benchmarks have been developed [Mialon2023, Rein2025]. Of particular interest in such evaluation scenarios is to identify setups (agent scaffolds [Yao2022]) that increase the success probability of models, and to make robust predictions for future behavior in deployment contexts. This extends to, for example, which tools LLMs have at their disposal, or under which level of reasoning effort they operate. The assessment of these effects has far-reaching financial implications, as token costs for advanced reasoning LLMs easily reach hundreds of US Dollars per task.
@@
Investigating the raw data per task, we observed an asymmetric bimodal success rate distribution across tasks (see Figure 7). This is a common finding in agentic evaluation settings – LLMs either pass or fail at all their attempts (success rate either 0 or 1) for most tasks, an intermediate level of success is rare. This suggests that such evaluation items are either saturated, or that they are too difficult, and therefore do not provide fine-grained diagnostic value to meaningfully differentiate LLM capabilities. Binomial GLMs struggle with over-dispersed bimodal outcome distributions (high prevalence of 0s and 1s), as success rates are assumed to be drawn from unimodal distributions. Too account for such bimodal outcome data distributions, we can instead use Beta-Binomial GLMs, which assume that each of the binomial outcome observations has an individual success probability.
@@
Figure 9: Forest plot for the effect of reasoning effort (none, low, intermediate, high) and task difficulty (easy, mid, hard) on performance of 4 different LLMs (use case 3). The plot displays posterior means, posterior distributions and 95% HPDI (on logit scale) and provides an indication for the width (thin lines) and uncertainty of the posterior parameter distribution (thick lines, 95% HPDI), around the posterior mean (blue dots) for the respective parameter estimates.

### v3

@@ 2.1 Multilevel (hierarchical) linear modeling using Generalized Linear Models
$\begin{split}Y_{i}&\sim\text{Binomial}(1,p_{i})\\ \text{logit}(p_{i})&=\beta_{0}+\beta_{1}X_{1i}+\cdots+\beta_{k}X_{ki}\end{split}$ (1)
where $Y_{i}$ denotes the number of successes in a binomial distribution. In the specific case where $n=1$, $Y_{i}$ corresponds to a Bernoulli trial, resulting in a binary outcome. The success probability $p_{i}$ is linked to the linear predictor via the logit link function.
@@ 3.1 Practical considerations for fitting GLMs using HiBayES
    • The number of effective samples (n_eff) is a statistical estimate of how many independent samples the dependent samples are equivalent to after accounting for autocorrelation. This number gives an indication of how well-defined the GLM is and how smoothly the sampling process proceeded. This should be $\geq$ the number of MCMC samples drawn per chain (e.g., 4000). In general, a higher n_eff is indicative of a well-specified and efficient model that produces efficient sequential samples.
@@
Figure 3: Highest posterior density intervals (HPDIs) used for hypothesis testing. Depicted are posterior probability distributions estimating two LLMs’ average performance and 95% HPDIs of these posterior distributions (both on probability scale) from simulated data fitted with a GLM. The HPDIs represent the interval of parameter values that is most consistent with the data – the range of parameter values that the true performance score is most likely to fall into. HPDIs in panel A overlap completely (>99% overlap), which suggests that LLM performance is equivalent (LLM A == LLM B). In panel B, HPDIs overlap partially ($\sim$88overlap), hence the result is inconclusive (LLM B <> LLM A). In panel C, HPDIs do not overlap (0% overlap), so we can conclude that LLMs perform differently (LLM B > LLM A).
@@ 3.3 Use case 1: One LLM evaluated on multiple domains (benchmarks) with varying numbers of items
Figure 5: Forest plot to display posterior means, posterior distributions, and 95% HPDI (on probability scale) for use case 1. Focusing on domain-specific and overall performance effects (blue), this plot indicates the uncertainty (vertical lines, 95% HPDI), around the posterior mean (dots) for the respective parameter estimates. Additionally, for comparison between HiBayES and conventional evaluation statistics approaches, we present empirical means and SEM per domain (orange) and for overall (red).
While the HiBayES GLM estimated domain-specific effects and the empirical means per domain are similar, the empirical SEMs are much narrower than the GLM estimated HPDIs. This may be misleading and produce overly confident interpretations of LLM capabilities, as the empirical SEM provides too narrow estimates of uncertainty. However, an even larger risk of overstated capability interpretations stems from the overestimation of overall performance levels using the empirical mean. Here, the estimate is influenced strongly by the larger number of datapoints from domain 2 ($N=6768$) vs domain 1 ($N=1257$). While this could potentially be mitigated by averaging per domain first and then computing the overall performance, such an approach will not be reliable for estimating the SEM. Using HiBayES the overall performance is estimated exactly at the central point between Coding and Reasoning estimates (as one would intuitively expect) and the HPDI spans the width of both domain-specific HPDIs – the overall mean estimate regularizes the domain-specific estimates.
@@
$\text{logit}(p_{i,j,k})$ $=\text{subdomain}_{i,j,k}$
$\text{model}_{i}$ $=\mu_{\text{overall}}+\sigma_{\text{model}}\cdot z_{\text{model}_{i}}$
@@
$\sigma_{\text{model}}$ $\sim\text{HalfNormal}(0.1)$
$z_{\text{model}_{i}}$ $\sim\text{Normal}(0,1)$
$\text{domain}_{i,j}$ $=\text{model}_{i}+\sigma_{\text{domain}}\cdot z_{\text{domain}_{i,j}}$
@@
$z_{\text{domain}_{i,j}}$ $\sim\text{Normal}(0,1)$
$\text{subdomain}_{i,j,k}$ $=\text{domain}_{i,j}+\sigma_{\text{subdomain}}\cdot z_{\text{subdomain}_{i,j,k}}$
$\sigma_{\text{subdomain}}$ $\sim\text{HalfNormal}(0.1)$
$z_{\text{subdomain}_{i,j,k}}$ $\sim\text{Normal}(0,1)$
@@
$\text{Correct}_{i,j,k}$ denotes the number of correct answers for the $i$-th LLM, $j$-th domain, and $k$-th subdomain, distributed as the proportion of correct answers (proportion parameter $p$) among the number of trials ($n_{i,j,k}$). The GLM assumes that the individual LLM-specific parameters are drawn from an LLM distribution. The individual domain-specific parameters are drawn from a domain distribution that is parameterized by the LLM-specific effects. The individual subdomain parameters are drawn from a subdomain distribution that is parameterized by the domain-specific effects.
@@
Figure 6 shows the posterior parameter estimates and compares estimated model (LLM), domain and subdomain effects and the uncertainty of the estimates (HPDI) to empirical means (arithmetic means) and standard errors of the mean (SEM) for each LLM, domain and subdomain separately (Figure 6). We observe that the empirical SEMs are consistently narrower than the GLM estimated HPDIs. Furthermore, the empirical means and GLM estimated parameter means are similar, but not identical. The latter phenomenon, called “shrinkage,” refers to partial pooling of parameters across nested data levels in hierarchical GLMs and a resulting contraction of lower-level parameter estimates (e.g., subdomain means) towards higher-level estimates (e.g., domain means). This form of regularization imposed by the hierarchical data structure helps to produce less over-fitting and more stable estimates. In other words, the parameter estimates derived from hierarchical GLMs allow us to more confidently generalize the capability in question to new contexts beyond the specific evaluation – e.g., real-world deployment contexts.
@@
Figure 6: Forest plot to display posterior means and 95% HPDI (on the probability scale) for use case 2. Focusing on model (LLM)-, domain- and subdomain-specific effects, this plot indicates the uncertainty (vertical lines, 95% HPDI), around the posterior mean (dots) for the respective parameter estimates. Additionally, for comparison between HiBayES and conventional evaluation statistics approaches, we present empirical means and SEM per LLM, domain, subdomain (orange).
The HPDIs for the subdomains estimated in a hierarchical GLM suggest similar differences found using the conventional approach (95%-HPDI DS-1000 = [.34;.40].34.40[.34;.40][ .34 ; .40 ] vs [.47;.53].47.53[.47;.53][ .47 ; .53 ] and 95%-HPDI BoolQ = [.75;.78].75.78[.75;.78][ .75 ; .78 ] vs [.80;.83].80.83[.80;.83][ .80 ; .83 ]). Additionally, HPDIs allow to quantify the magnitude and uncertainty around these differences in a more nuanced and continuous way – supporting an inference that the performance difference observed for BoolQ is smaller than for DS-1000. The conventional approach only allows us to make the discrete inference that LLM performance levels differ significantly on both subdomains.
@@
An additional advantage of HiBayES is that GLMs like the one shown in Equation 3 straightforwardly extend to evaluation scenarios that cover more than two domains or more than two LLMs (e.g., comparing 3 LLMs) without adjusting the statistical inference process, correcting for multiple comparisons, or using another family of tests. Since all estimated parameters are indexed, it is possible to add another number/index and estimate 3 (x3) LLM/domain parameters instead of 2 (x2) – we illustrate this below in use case 3.
@@
Interest in how AI models behave in real-world, agentic settings has surged in recent years. To test these capabilities, more complex, challenging and extended benchmarks than QA benchmarks have been developed [Mialon2023, Rein2025]. Of particular interest in such evaluation scenarios is to identify setups (agent scaffolds [Yao2022]) that increase the success probability of models, and to make robust predictions for future behavior in deployment contexts. This extends to, for example, which tools LLMs have at their disposal, or under which level of reasoning effort they operate. The assessment of these effects has far-reaching financial implications, as token costs for advanced reasoning LLMs easily reach hundreds of US Dollars per task.
@@
Investigating the raw data per task, we observed an asymmetric bimodal success rate distribution across tasks (see Figure 7). This is a common finding in agentic evaluation settings – LLMs either pass or fail at all their attempts (success rate either 0 or 1) for most tasks, an intermediate level of success is rare. This suggests that such evaluation items are either saturated, or that they are too difficult, and therefore do not provide fine-grained diagnostic value to meaningfully differentiate LLM capabilities. Binomial GLMs struggle with over-dispersed bimodal outcome distributions (high prevalence of 0s and 1s), as success rates are assumed to be drawn from unimodal distributions. To account for such bimodal outcome data distributions, we can instead use Beta-Binomial GLMs, which assume that each task has its own task-specific success rate, drawn from a distribution shared across tasks.
@@
Figure 9: Forest plot for the effect of reasoning effort (none, low, intermediate, high) and task difficulty (easy, mid, hard) on performance of 4 different LLMs (use case 3). The plot displays posterior means and 95% HPDI (vertical lines) for the respective parameter estimates (on logit scale).